Patriot 12 meter telescope

Last Update: 28jun22

Links
Taking data with the 12meter
History:
acceptance test
Measurements
debugging the hardware
cal values
rfi measurements
    rfi movies
Specs

Monitor tsys with hardware winking cal data

Measured power Levels
xband receiver chain levels
System performance
The telescope model
Documentation
my notes
fieldsystem, rdbe, mark6 page
Looking at the sun
New wide band receiver

Pulsars observing with the12 meter


Daily temperature plots from 12meter hub,ctrlbox, and weatherStation.

Pictures of the 12 meter system
    receiver chain

12meter control program


Trouble shooting polA 210920

Debugging history:


History: (top)


All Measurements:




Specs:


az
el


position limits
-180 to +360 deg
5.8 to 88 deg



sband
xband
date
notes
freqRange:
2.185 - 2.364 GHz
8.05 - 9.2 GHz

rf filters @12m feb16

2.208 - 2.343 GHz

11mar22
new sband rf filter installed





bmWdith
fwhm
40 Amin
10 Amin
mar22

secondary apex to
feedhorn mouth
18.125" (46cm)


measured 26may22

SPD @ fiber
rcvr output
-81.5 dbm polA
-76.0 dbm polB
-76.4 dbm polA
-70.5 dbm polB
160328
  • 3Mhz rbw, average 100.
  • 1db extra pad in xband polB (10mar16)
  • switch tx,rx so each pair is same serial #
  • 5ft cable rcvr to specAna.
  •  .5db loss/ft at 8.6ghz
  • signal 2.5 db higher at xb at rcver.
  • polA has 5 to 6 db less power then polB

UDC output levels 500-1000 vlbi outputs: dbm/Hz
skyCfr (Ghz)
polA
polB


2.19
2.27
2.364
-94.1
-94.7
-97.2
-90.4
-88.7
-93.5
160331
udc atten was set to 15.5 db atten
8.1
8.6
9.1
-89.5
-89.9
-91.4
-85.1
-84.5
-85.3


Schematic of system.


Fiber optic output (sband not terminated)

 Measurements with power meter.

spectra fiber optic output

Fiber tx,rx performance

Equivalent noise Temperature at FO input/output for sband





udc levels



Xband power levels

210426 - power meter in pedestal and cone

210621: lna output looking at blank sky

    On 11jun21 felix hooked the field fox spectrum analzyer to the output of the lna's while looking at blank sky:

The plots show the measured  spectrum output from the lnas (.ps) (.pdf)

processing: x101/210611/12m_lnaoutput.pro


sband power levels

210426 power meter  in pedestal

220311 power meter in pedestal after new sband filter installed


xband receiver chain levels

    The parameters of the components in the xband receiver chain were used to compute the levels as you move through the chain.

Notes:

input file format

    The input file (for the original configuration) was:

# 12meter xband chain
# if possible use values 8-9ghz
#name  noiseFigure/Tsys   gain/loss db
#
sky           10K                0db
lna           41K                40db   from patriot measurements
cLnaToAmp2   -                 -10db    26apr21 spectrum analzyer measurements  8-9 ghz
hubAmp2       5db                20db   26apr21  measurement 8-9 ghz. datasheet says 18
cAmp2ToPed   -                 -7.5db   26apr21 total power measurements
amp3PostA   7db                24db     datasheet
padr3          -                 -3db   dana drawing
xbfilter     -                   -2db   dana drawing
padr4          -                 -5db   dana drawing
amp4PostA   7db                24db     dana drawing
padr5          -                 -2db   data drawing
xbdiplexor     -                 -1db   dana drawing
padR20         -                 -7db   dana drawing

Key:


Xband receiver chain levels
date
link
state
description
210621
(.txt)

sys config before xb filter moved
210622
(.txt)
after xb filter moved before 1st amp in postamp chassis
210623
(.txt)
remove hub amp for testing (testing). will eventually  replace them
210623
(.txt) not used
remove hub amp, replace 5db pad with 3db pad (testing)
210629
(.txt)
current
polB 20db hub amp bad. temporarily replace both amps with  11.5 db amps

Notes:

sband receiver chain levels


date
link
state
description
220310
(.txt)
testing
test values.







Documentation: (top)


Daily temperature plots

    Daily plots are made of the temperature in the 12meter hub, 12meter motor control box, and the temperatures measured from the orion weather station (atop the lidar lab). The plots are generated at 1am each day via  a crontab  running on megs3. A postcript (.ps) and pdf (.pdf) file are generated for each day.

    Directory listing of daily temperature plots


processing: x101/12meter/temp/plottempdaily.sc



My notes


History extended


210409 - fiber patch cords in pedestal replaced

    When opening and closing the cabinet (holding postamps,fiber tx) in the pedestal , saw large change in the levels in the ctrlroom (1/6 Tsys). We found the polB fiber connector at the patch  panel loose. When we tightened it, the output power in the ctrlroom went  way down. When loosening it, it went back to the original level.
    Osvaldo changed both of the fiber patch cables. After  this, opening and closing the cabinet made little change in the power levels seen at  the control room. We still much larger dropouts in polB relative to polA (although they both seem to happen together).  May be some kind out of band compression? More notes: x101/210409/chksys.txt


<- page up
 home_~phil